Circular No. 7631

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001bs IN UGC 10018 S. Beckman and W. D. Li, University of California at Berkeley, report

the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on May 22.4 UT (mag ∼ 18.3) and May 23.3 UT (mag ∼ 18.2). The new object is located at α = 15^{h}43^{m}39^{s}.38, δ = +67^{o}45′31′′.3 (equinox 2000.0), which is 0′′.9 west and 7′′.7 south of UGC 10018. A KAIT image of the same field taken on May 7.4 UT showed nothing at this position (limiting mag ∼ 19.0).

SUNGRAZING COMETS Further to IAUC 7613, MPEC 2001-H44, 2001-J3, 2001-J35 and 2001-

K19 give detailed information on six more Kreutz sungrazing comets, the SOHO discovery observations of which are:

2001 UT  α_{2000}  δ_{2000}  Comet
Apr. 26.488  2^{h}29.^{m}2  +11^{o}49′  C/2001 H6
30.188  2 35.6  +13 37  C/2001 H7
May 5.354  2 55.1  +15 04  C/2001 J2
6.729  2 56.4  +15 05  C/2001 J3
10.131  3 12.8  +16 16  C/2001 J4
19.471  3 51.6  +17 35  C/2001 K2

COMET C/2001 K3 (SKIFF) B. A. Skiff, Lowell Observatory, reports his discovery of a comet on

images taken with the LONEOS 0.59-m Schmidt on May 22.4 (discovery position below); the coma was well condensed, ∼ 20′′ in diameter, and a tail extended ∼ 60′′ in p.a. 225^{o}. Following placement in The NEO Confirmation Page, further observations were reported, and they are listed on MPEC 2001-K24, together with preliminary parabolic orbital elements (T = 2001 Jan. 12, q = 1.87 AU, ω = 315^{o}, Ω = 281^{o}, i = 37^{o}, equinox 2000.0). M. Tichý and M. Kočer (Kleť, 0.57-m reflector) reported a compact 8′′ coma; D. T. Durig (Sewanee, TN, 0.30-m reflector) a 30′′ tail; K. Smalley (Olathe, KS, 0.75-m reflector) a tail ∼ 30^{o} wide, brightest along the southern edge (p.a. ∼ 210^{o}), where it extended for ∼ 2′; R. Dyvig (Quinn, SD, 0.66-m reflector) a possible coma and faint tail in p.a. ∼ 225^{o}.

2001 UT  α_{2000}  δ_{2000}  m_{1}
May 22.39643  20^{h}59^{m}32.^{s}00  −6^{o}28′00′′.2  16.0
2001 May 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Brian G. Marsden